# Multi-Layer Cycle Benchmarking for high-accuracy error characterization

### v1

@@ Appendix B Equivalence between products of fidelities
The objective of this section is to establish the equivalence between products of fidelities directly measurable via Multi-Layer Cycle Benchmarking (MLCB)—such as $o_{3},o_{4},o^{\prime}_{4},o_{5},c_{4}$ introduced in the main text—and specific ratios of unlearnable fidelities, $\mu_{q}^{(BG)}$. As established at the end of Sec. IV.1, two functions of fidelities are deemed equivalent if one can be fully derived from the other using only high-accuracy data, i.e. (products of) learnable fidelities.
@@
To achieve this, we first introduce an equivalence test to formally prove the equivalence between functions of fidelities within a single layer. We then extend this test to multi-layer scenarios and demonstrate its application using a concrete example of three-qubit chains. This example illustrates each step required to prove equivalence and establish relationships such as those expressed in Eqs. (15)–(17). Building on this example, we generalize the methodology to more complex cases and summarize the results.
B.1 Equivalence test
@@
We begin by focusing on a single layer $C$ and linear combinations of logarithms of fidelities. These combinations can be written as $\zeta_{Y}=\sum_{\alpha\in\mathcal{X}_{Y}}\gamma_{\alpha}\log f^{C}_{\alpha}$, where $\gamma^{Y}_{\alpha}$ are numerical coefficients and $\mathcal{X}_{Y}$ is a set of Pauli strings.
@@
Within the Sparse Pauli-Lindblad (SPL) model, the vector of model parameters $\vec{\lambda}$ determines all functions of fidelities via Eq. (8). In particular, the vector $\vec{\Xi}$, consisting of all learnable (products of) fidelities which can be determined, for instance, by following the strategy outlined in Section II.2, is determined by
@@
To test whether two functions of fidelities, $\zeta_{A}$ and $\zeta_{B}$, are equivalent, we compute $\bar{M}$, $\vec{m}_{A}$ and $\vec{m}_{B}$. Let $r$ be the rank of $\bar{M}$; $r_{Y}$ the rank of matrix $M$ concatenated with the extra row $(\vec{m}_{Y})^{T}$; $r_{AB}$ the rank of the matrix $\bar{M}$ concatenated with $(\vec{m}_{A})^{T}$ and $(\vec{m}_{B})^{T}$. The equivalence condition is then
@@
Figure 9: Visualization of single-row matrices $(\vec{m}_{\alpha})^{T}$ associated with the blue layer (panel a) and the green layer (panel b) for the three-qubit open chain depicted in Fig. 3(c). On the horizontal axes, all the $27$ ordered Pauli strings $\kappa\in\mathcal{K}$ are listed. The first nine consists are weight-one Pauli strings, the second nine are weight-two Pauli strings with support on qubits 00 and $1$, and the last nine are weight-two Pauli strings with support on qubits $1$ and $2$. As for the Pauli strings $\alpha$, they are specified to the left of each row. Note that the last row actually depicts $(\vec{m}_{j}^{T})=(\vec{m}_{IIX}^{T})+(\vec{m}_{IZX}^{T})$.
B.2 Three-qubit open chain
We now consider the three-qubit open chain described in Sec. IV.1. Our goals are: (i) to prove the equivalence between $o_{3}=f^{B}_{XIX}f^{G}_{XZX}$ and $\mu_{1}^{(BG)}=f^{B}_{XII}/f^{G}_{IIX}$ (ii) to prove Eqs. (15-17) which directly link these quantities.
For concreteness, we assume a Quantum Processing Unit (QPU) with square topology, ensuring qubits 00 and $2$ are not neighbours. In this case, the SPL model describing the noise in the chain features a total of $|\mathcal{K}|=27$ parameters, associated with the Pauli strings $\kappa$ listed on the $x$-axes of the plots in Fig. 9.
We first analyze the blue layer, focusing on the function of (now trivial) fidelities $F^{B}_{1}=f^{B}_{XII}$ and $F^{B}_{2}=f^{B}_{XIX}$. Computing the corresponding matrix rows
@@ B.2 Three-qubit open chain
we verify the equivalence condition from Eq. (28). This proves $\log(F^{B}_{1})$ can be expressed as a linear combination of $\log(F^{B}_{2})$ and other function of learnable fidelities. To explicitly identify the latter, we implement a simple search algorithm, which progressively attempt to get rid of (randomly chosen) rows of matrix $\bar{M}$ while still preserving the condition on the ranks in Eq. (28). We find that, in this case, the single row of matrix $\bar{M}$ associated with the learnable fidelity $f^{B}_{IIX}$, i.e.
@@
with $L=B$, $\epsilon^{B}=1$, $\vec{\Sigma}^{B}=(-1)$ and $\vec{F}^{B}=(f^{B}_{IIX})$. This proves the validity of 15.
@@
where the last row $(\vec{m}_{j}^{T})=(\vec{m}_{IIX}^{T})+(\vec{m}_{IZX}^{T})$ is associated to the learnable product of fidelities in the orbit $\mathcal{R}^{G}_{IIX}=\{P_{IIX},P_{IZX}\}$. Such a relation is visualized in Fig. 9(b). This result can be summarized in the form of Eq. (32), with $L=G$, $\epsilon^{G}=-1$, $\vec{\Sigma}^{G}=(1,-1,1)$ and $\vec{F}^{G}=(f^{G}_{XZI},f^{G}_{IZI},f^{G}_{IIX}f^{G}_{IZX})$. This proves the validity of 16.
By combining these results, valid for each layer individually, we obtain Eq. 17, thus showing the equivalence between $o_{3}=F^{B}_{2}F^{G}_{2}$ and $\mu_{1}^{(BG)}=F^{B}_{1}/F^{G}_{1}$. It is important to stress that the equivalence class does not consists only of these two functions of fidelities. It is indeed possible to find other equivalent functions of fidelities, all associated to the same Degrees of Freedom (DOF), whose learnability is unlocked by MLCB. For example, one has
@@ B.3 Other chains
Longer chains introduce additional challenges due to the connectivity of the underlying QPU. Non-neighboring qubits in the chain may still interact via the hardware’s topology, increasing the number of SPL parameters and altering the relationships between fidelity functions. Relevant examples are shown in the first column of Tables 2-5
@@
In the following, we analyze all relevant chain configurations, considering a QPU with square topology. For each case, we apply the methodology outlined for the three-qubit chain. Specifically, focusing on one layer at a time, we use the equivalence test to demonstrate the equivalence between two functions of fidelities, denoted as $F^{L}_{1}$ and $F^{L}_{2}$, respectively. This approach enables us to prove the equivalence—and derive a direct connection—between the fidelity product measurable via MLCB, $F^{B}_{2}F^{G}_{2}$, and the ratios of unlearnable fidelities, $F^{B}_{1}/F^{G}_{1}$. The results for all chain configurations are summarized in Tables 2-5.
@@
It is important to note that the reported $\vec{\Sigma}^{L}$ and $\vec{F}^{L}$ are not unique, as alternative choices for these quantities can still satisfy Eq. (32). An interesting open question concerns the optimization of the choice of $\vec{\Sigma}^{L}$ and $\vec{F}^{L}$, with the goal of minimizing the number of components in each vector for a given chain. While this optimization does not affect the learnability of fidelities and is thus beyond the scope of this paper, it can be beneficial when implementing the characterization procedure in practice. For instance, in the four-qubit closed chain analyzed in Table 5, connecting $c_{4}=f^{B}_{IXXX}f^{G}_{ZXYY}f^{B}_{IYYX}f^{G}_{ZYXY}$ to the simple fidelity ratio $\mu_{q}^{(BG)}$ requires considering a total of $28$ learnable (products of) fidelities. Reducing this number could streamline the analyses and make them more robust with respect to the spread of statistical uncertainties.
Appendix C Fitting the noise parameters
@@
Here we detail how we impose the high-accuracy constraint obtained with MLCB on the low-accuracy fidelities. Let us focus on a qubit $q$, belonging to the support of $l$ layers $L_{1},\dots,L_{l}$. When relying exclusively on conventional CB protocols, qubit $q$ is associated with $l$ unlearnable DOF, which can be identified with the $l$ fidelities in $\{f^{L_{j}}_{\beta_{q}^{j}}\}$, where the weight-one Pauli strings $\beta_{q}^{j}$ feature a single $X$ operator on the qubit connected to $q$ via the $L_{j}$ layer. Low-accuracy methods, like the symmetry assumption or unit-depth CB-like experiments, provide us with estimates for each one of these fidelities, which we indicate with an extra tilde $\tilde{f}^{L_{j}}_{\beta_{q}^{j}}$ and represent the numerical values used for the conventional model fitting.
@@
MLCB allows us to determine $l-1$ additional high-accuracy constraints, which can be identified, for example, with fidelity ratios
@@
for $j=1,\dots,l-1$. This reduces the number of unlearnable DOF to one, which we can identify with a specific fidelity, say $f_{q}^{L_{1}}$. We estimate its numerical value $u$ by minimizing the sum $S$ of the squared residues
@@
computed by comparing the low-accuracy fidelity estimations with the effect of the high-fidelity constraints from MLCB. This procedure can be readily adjusted to tackle scenarios where different pairs of layers have been studied with MLCB.
Appendix D Random noise models
@@
Inactive parameters are drawn from a Gaussian distribution with mean $m_{w}^{i}$ and standard deviation $\sigma_{w}^{i}$. The active ones are sampled from a Gaussian with mean $m_{w}^{a}$ and standard deviation $\sigma_{w}^{a}$, where, for every gate in the layer, $m_{w}^{a}$ itself is drawn from a Gaussian with mean $\bar{m}_{w}^{a}$ and standard deviation $\bar{\sigma}_{w}^{a}$ This additional sampling step introduces variability between gates within a single layer, resulting in better and worse-performing ones.
@@
[…]
@@
[…]
@@
$f^{B}_{XIII}$ $f^{B}_{XIXI}f^{B}_{XIXZ}$ $+0.50$ $\begin{pmatrix}-0.50\\ \end{pmatrix}$ $\begin{pmatrix}f^{B}_{IIXZ}f^{B}_{IIXI}\\ \end{pmatrix}$
@@
$f^{G}_{IIXI}$ $f^{G}_{XZXZ}f^{G}_{XZXI}$ $-0.50$ $\begin{pmatrix}+0.50\\ -0.50\\ +0.25\\ +0.75\\ \end{pmatrix}$ $\begin{pmatrix}f^{G}_{XZII}f^{G}_{XZII}\\ f^{G}_{IZII}f^{G}_{IZII}\\ f^{G}_{IIXZ}f^{G}_{IZXZ}\\ f^{G}_{IZXI}f^{G}_{IIXI}\\ \end{pmatrix}$
@@
$f^{B}_{XIII}$ $f^{B}_{XIXI}f^{B}_{XIXZ}$ $+0.50$ $\begin{pmatrix}+0.25\\ -0.50\\ +0.25\\ -0.25\\ \end{pmatrix}$ $\begin{pmatrix}f^{B}_{IIIZ}f^{B}_{IIIZ}\\ f^{B}_{IIXZ}f^{B}_{IIXI}\\ f^{B}_{XZII}f^{B}_{XIII}\\ f^{B}_{XIIZ}f^{B}_{XZIZ}\\ \end{pmatrix}$
@@
$f^{G}_{IIXI}$ $f^{G}_{XZXZ}f^{G}_{XZXI}$ $-0.50$ $\begin{pmatrix}-0.50\\ +0.25\\ +0.75\\ -0.25\\ +0.50\\ -0.25\\ +0.25\\ \end{pmatrix}$ $\begin{pmatrix}f^{G}_{IZII}f^{G}_{IZII}\\ f^{G}_{IIXZ}f^{G}_{IZXZ}\\ f^{G}_{IIXI}f^{G}_{IZXI}\\ f^{G}_{IIIZ}f^{G}_{IIIZ}\\ f^{G}_{XZII}f^{G}_{XZII}\\ f^{G}_{XIII}f^{G}_{XIII}\\ f^{G}_{XIIZ}f^{G}_{XIIZ}\\ \end{pmatrix}$
@@
$f^{B}_{IXII}$ $f^{B}_{IXIX}f^{B}_{ZXIX}$ $+0.50$ $\begin{pmatrix}-0.25\\ +0.25\\ -0.50\\ \end{pmatrix}$ $\begin{pmatrix}f^{B}_{ZXII}f^{B}_{ZXZI}\\ f^{B}_{IXZI}f^{B}_{IXII}\\ f^{B}_{IIIX}f^{B}_{IIIX}\\ \end{pmatrix}$
@@
$f^{G}_{IIXI}$ $f^{G}_{XZXZ}f^{G}_{XZXI}$ $-0.50$ $\begin{pmatrix}-0.50\\ +0.50\\ +1.00\\ \end{pmatrix}$ $\begin{pmatrix}f^{G}_{IIZI}f^{G}_{IIZI}\\ f^{G}_{IXZI}f^{G}_{ZXZI}\\ f^{G}_{IIZX}f^{G}_{IIIX}\\ \end{pmatrix}$
@@
$f^{B}_{IXII}$ $f^{B}_{IXIX}f^{B}_{ZXIX}$ $+0.50$ $\begin{pmatrix}+0.25\\ -0.25\\ -0.25\\ +0.25\\ -0.25\\ \end{pmatrix}$ $\begin{pmatrix}f^{B}_{ZIII}f^{B}_{ZIII}\\ f^{B}_{IIIX}f^{B}_{IIIX}\\ f^{B}_{ZXII}f^{B}_{ZXZI}\\ f^{B}_{IXZI}f^{B}_{IXII}\\ f^{B}_{ZIIX}f^{B}_{ZIIX}\\ \end{pmatrix}$
@@
$f^{G}_{IIXI}$ $f^{G}_{XZXZ}f^{G}_{XZXI}$ $-0.50$ $\begin{pmatrix}-0.25\\ +0.25\\ +0.50\\ -0.50\\ +0.75\\ \end{pmatrix}$ $\begin{pmatrix}f^{G}_{ZIII}f^{G}_{ZIII}\\ f^{G}_{ZIIX}f^{G}_{ZIZX}\\ f^{G}_{IXZI}f^{G}_{ZXZI}\\ f^{G}_{IIZI}f^{G}_{IIZI}\\ f^{G}_{IIZX}f^{G}_{IIIX}\\ \end{pmatrix}$

### v2

@@ Appendix C Equivalence between products of eigenvalues
The objective of this section is to establish the equivalence between products of eigenvalues directly measurable via Multi-Layer Cycle Benchmarking (MLCB)—such as $o_{3},o_{4},o^{\prime}_{4},o_{5},c_{4}$ introduced in the main text—and specific ratios of unlearnable eigenvalues, $\mu_{q}^{(BG)}$. As established at the end of Sec. IV.1, two functions of eigenvalues are deemed equivalent if one can be fully derived from the other using only high-accuracy data, i.e. (products of) learnable eigenvalues.
@@
To achieve this, we first introduce an equivalence test to formally prove the equivalence between functions of eigenvalues within a single layer. We then extend this test to multi-layer scenarios and demonstrate its application using a concrete example of three-qubit chains. This example illustrates each step required to prove equivalence and establish relationships such as those expressed in Eq.(15). Building on this example, we generalize the methodology to more complex cases and summarize the results.
C.1 Equivalence test
@@
We begin by focusing on a single layer $C$ and linear combinations of logarithms of eigenvalues. These combinations can be written as $\zeta_{Y}=\sum_{\alpha\in\mathcal{X}_{Y}}\gamma_{\alpha}\log f^{C}_{\alpha}$, where $\gamma^{Y}_{\alpha}$ are numerical coefficients and $\mathcal{X}_{Y}$ is a set of Pauli strings.
@@
Within the Sparse Pauli-Lindblad (SPL) model, the vector of model parameters $\vec{\lambda}$ determines all functions of eigenvalues via Eq. (6). In particular, the vector $\vec{\Xi}$, consisting of all learnable (products of) eigenvalues which can be determined, for instance, by following the strategy outlined in Section III.1, is determined by
@@
To test whether two functions of eigenvalues, $\zeta_{A}$ and $\zeta_{B}$, are equivalent, we compute $\bar{M}$, $\vec{m}_{A}$ and $\vec{m}_{B}$. Let $r$ be the rank of $\bar{M}$; $r_{Y}$ the rank of matrix $M$ concatenated with the extra row $(\vec{m}_{Y})^{T}$; $r_{AB}$ the rank of the matrix $\bar{M}$ concatenated with $(\vec{m}_{A})^{T}$ and $(\vec{m}_{B})^{T}$. The equivalence condition is then
@@
Figure 9: Visualization of single-row matrices $(\vec{m}_{\alpha})^{T}$ associated with the blue layer (panel a) and the green layer (panel b) for the three-qubit open chain depicted in Fig. 3(c). On the horizontal axes, all the $27$ ordered Pauli strings $\kappa\in\mathcal{K}$ are listed. The first nine consists are weight-one Pauli strings, the second nine are weight-two Pauli strings with support on qubits 0 and $1$, and the last nine are weight-two Pauli strings with support on qubits $1$ and $2$. As for the Pauli strings $\alpha$, they are specified to the left of each row. Note that the last row actually depicts $(\vec{m}_{j}^{T})=(\vec{m}_{IIX}^{T})+(\vec{m}_{IZX}^{T})$.
C.2 Three-qubit open chain
We now consider the three-qubit open chain described in Sec. IV.1. Our goals are: (i) to prove the equivalence between $o_{3}=f^{B}_{XIX}f^{G}_{XZX}$ and $\mu_{1}^{(BG)}=f^{B}_{XII}/f^{G}_{IIX}$ (ii) to prove Eq. (15) which directly links these quantities.
For concreteness, we assume a Quantum Processing Unit (QPU) with square topology, ensuring qubits 0 and $2$ are not neighbours. In this case, the SPL model describing the noise in the chain features a total of $|\mathcal{K}|=27$ parameters, associated with the Pauli strings $\kappa$ listed on the $x$-axes of the plots in Fig. 9.
We first analyze the blue layer, focusing on the function of (now trivial) eigenvalues $F^{B}_{1}=f^{B}_{XII}$ and $F^{B}_{2}=f^{B}_{XIX}$. Computing the corresponding matrix rows
@@ C.2 Three-qubit open chain
we verify the equivalence condition from Eq. (29). This proves $\log(F^{B}_{1})$ can be expressed as a linear combination of $\log(F^{B}_{2})$ and other function of learnable eigenvalues. To explicitly identify the latter, we implement a simple search algorithm, which progressively attempts to remove (randomly chosen) rows of matrix $\bar{M}$ while still preserving the condition on the ranks in Eq. (29). We find that, in this case, the single row of matrix $\bar{M}$ associated with the learnable eigenvalue $f^{B}_{IIX}$, i.e.
@@
with $L=B$, $\epsilon^{B}=1$, $\vec{\Sigma}^{B}=(-1)$ and $\vec{F}^{B}=(f^{B}_{IIX})$. This proves
$f^{B}_{XII}=\frac{f^{B}_{XIX}}{f^{B}_{IIX}}$ (34)
@@
where the last row $(\vec{m}_{j}^{T})=(\vec{m}_{IIX}^{T})+(\vec{m}_{IZX}^{T})$ is associated to the learnable product of eigenvalues in the orbit $\mathcal{R}^{G}_{IIX}=\{P_{IIX},P_{IZX}\}$. Such a relation is visualized in Fig. 9(b). This result can be summarized in the form of Eq. (33), with $L=G$, $\epsilon^{G}=-1$, $\vec{\Sigma}^{G}=(1,-1,1)$ and $\vec{F}^{G}=(f^{G}_{XZI},f^{G}_{IZI},f^{G}_{IIX}f^{G}_{IZX})$. This proves
$f^{G}_{IIX}=\frac{f^{G}_{XZI}(f^{G}_{IZX}f^{G}_{IIX})}{f^{G}_{XZX}f^{G}_{IZI}}$ (36)
By combining these results, valid for each layer individually, we obtain Eq. 15, thus showing the equivalence between $o_{3}=F^{B}_{2}F^{G}_{2}$ and $\mu_{1}^{(BG)}=F^{B}_{1}/F^{G}_{1}$. It is important to stress that the equivalence class does not consists only of these two functions of eigenvalues. It is indeed possible to find other equivalent functions of eigenvalues, all associated to the same Degrees of Freedom (DOF), whose learnability is unlocked by MLCB. For example, one has
@@ C.3 Other chains
Longer chains introduce additional challenges due to the connectivity of the underlying QPU. Non-neighboring qubits in the chain may still interact via the hardware’s topology, increasing the number of SPL parameters and altering the relationships between functions of eigenvalues. Relevant examples are shown in the first column of Tables 2-5
@@
In the following, we analyze all relevant chain configurations, considering a QPU with square topology. For each case, we apply the methodology outlined for the three-qubit chain. Specifically, focusing on one layer at a time, we use the equivalence test to demonstrate the equivalence between two functions of eigenvalues, denoted as $F^{L}_{1}$ and $F^{L}_{2}$, respectively. This approach enables us to prove the equivalence—and derive a direct connection—between the eigenvalue product measurable via MLCB, $F^{B}_{2}F^{G}_{2}$, and the ratios of unlearnable eigenvalues, $F^{B}_{1}/F^{G}_{1}$. The results for all chain configurations are summarized in Tables 2-5.
@@
It is important to note that the reported $\vec{\Sigma}^{L}$ and $\vec{F}^{L}$ are not unique, as alternative choices for these quantities can still satisfy Eq. (33). An interesting open question concerns the optimization of the choice of $\vec{\Sigma}^{L}$ and $\vec{F}^{L}$, with the goal of minimizing the number of components in each vector for a given chain. While this optimization does not affect the learnability of eigenvalues and is thus beyond the scope of this paper, it can be beneficial when implementing the characterization procedure in practice. For instance, in the four-qubit closed chain analyzed in Table 5, connecting $c_{4}=f^{B}_{IXXX}f^{G}_{ZXYY}f^{B}_{IYYX}f^{G}_{ZYXY}$ to the simple eigenvalue ratio $\mu_{q}^{(BG)}$ requires considering a total of $28$ learnable (products of) eigenvalues. Reducing this number could streamline the analyses and make them more robust with respect to the spread of statistical uncertainties.
Appendix D Experimental details
In this appendix, we provide further details on the experimental execution and data analysis of the MLCB protocol discussed in Sec. V.
For each of the six pairs of layers $(L_{1},L_{2})$ depicted in Fig. 5(a), we perform a complete MLCB experiment. Each experiment consists of running circuits for five different sequence depths $d\in{2,4,8,16,32}$. For each depth, we use a total of $4\times 10^{4}$ measurement shots, which are distributed evenly across 20 random Pauli twirling instances. This results in $2000$ shots per unique circuit configuration. To gather more statistics and develop a feeling of the time stability of the results, we repeat this entire procedure three times back-to-back for every layer pair.
The statistical uncertainty of our measurements is estimated using a bootstrapping procedure. For each of the three experimental runs, we generate 100 bootstrap samples by resampling with replacement from both the measurement shots and the 20 random twirling instances. The final reported values and confidence intervals for $\Delta_{q}^{L_{1}L_{2}}$ are derived from the combined distribution of these bootstrapped results, averaged across the three consecutive runs. In Fig. 10, we present the complete set of measured $\Delta_{q}^{L_{1}L_{2}}$ values for all six layer pairs, averaged over the three consecutive experimental runs.
To assess the normality of the bootstrapped distributions, we compute the skewness $s_{k}$ and excess kurtosis $k_{u}$ for each difference $\Delta^{L_{1}L_{2}}_{q}$. With the exception of $\Delta^{\rm OR}_{5}$, which exhibits large skewness $s_{k}=-1$ and excess kurtosis $k_{u}=2.5$ (as also evident also from the asymmetric whiskers in Fig. 10), all the other differences features small values within the ranges $0.18\leq s_{k}\leq 0.25$ and $0.14\leq k_{u}\leq 0.32$. This confirms the approximate normality of the bootstrapped distributions, thereby validating the use of the associated ratios $\Omega^{L_{1}L_{2}}_{q}$, between mean and standard deviations, to assess their statistical compatibility with zero.
As detailed in Fig. 11(a), the products of Pauli eigenvalues measured by MLCB deviates from $1$ by around $0.1$. By contrast, the absolute differences $|\Delta_{q}^{L_{1}L_{2}}|$ remain below $0.006$ (see Fig. 10 and Fig. 11(b)). While these deviations are statistically significant, as established above, their magnitude indicates that violations of the symmetry assumption between unlearnable quantities contribute only a few percent to the overall value of the Pauli eigenvalues.
Figure 10: Summary of all measured asymmetry values $\Delta_{q}^{L_{1}L_{2}}$. Each panel corresponds to a different pair of layers, as indicated by the title. The box plots show the distribution of $\Delta_{q}$ for each hot qubit $q$ in the chain, averaged over three consecutive MLCB experiments. The whiskers indicate the 95% confidence interval derived from bootstrapping.
Figure 11: (a) Distribution of all the “infidelities” $1-\mathcal{F}_{\alpha(q)}^{L_{1}L_{2}}$ (panel a) and of all the absolute differences $|\Delta_{q}^{L_{1}L_{2}}|$ (panel b) measured by the MLCB protocol, for all $q$ and all layer pairs.
Let us now illustrate with a detailed example what is accessible to MLCB and standard CB protocols in terms of products of Pauli eigenvalues. To this end, we focus on the layers pair BR and on qubit $q=11$, i.e. to the case analyzed in detail in Fig. 5(c). Here we have
$\mathcal{F}_{\alpha(11)}^{BR}=f_{IXIX}^{B}f_{IXZX}^{R}f_{ZXIX}^{B}f_{ZXZX}^{R},$ (38)
where the Pauli strings refers to qubits $9$, $10$, $11$ and $12$ (see Fig. 5(a)). As $\beta(11)=IXZX$, we also have
$\mathcal{F}_{\beta(11)}^{BR}=f_{IXZX}^{B}f_{ZXIX}^{R}f_{ZXZX}^{B}f_{IXIX}^{R}.$ (39)
Those two products can be measured with high-accuracy with MLCB but cannot be individually accessed via conventional CB. The latter could only be used to measure, by combining the analyses of the two layers, the product $f=\mathcal{F}_{\alpha(11)}^{BR}\mathcal{F}_{\beta(11)}^{BR}$ which reads
$\begin{split}f=&(f_{IXIX}^{B}f_{IXZX}^{B})(f_{ZXIX}^{B}f_{ZXZX}^{B})\\ &\quad(f_{IXIX}^{R}f_{ZXZX}^{R})(f_{IXZX}^{R}f_{ZXIX}^{R})\end{split}$ (40)
and all the products within round brackets are learnable with standard CB.
Appendix E Fitting the noise parameters
@@
Here we detail how we impose the high-accuracy constraint obtained with MLCB on the low-accuracy eigenvalues. Let us focus on a qubit $q$, belonging to the support of $l$ layers $L_{1},\dots,L_{l}$. When relying exclusively on conventional CB protocols, qubit $q$ is associated with $l$ unlearnable DOF, which can be identified with the $l$ eigenvalues in $\{f^{L_{j}}_{\beta_{q}^{j}}\}$, where the weight-one Pauli strings $\beta_{q}^{j}$ feature a single $X$ operator on the qubit connected to $q$ via the $L_{j}$ layer. Low-accuracy methods, like the symmetry assumption or unit-depth CB-like experiments, provide us with estimates for each one of these eigenvalues, which we indicate with an extra tilde $\tilde{f}^{L_{j}}_{\beta_{q}^{j}}$ and represent the numerical values used for the conventional model fitting.
@@
MLCB allows us to determine l1l-1 additional high-accuracy constraints, which can be identified, for example, with eigenvalue ratios
@@
for $j=1,\dots,l-1$. This reduces the number of unlearnable DOF to one, which we can identify with a specific eigenvalue, say $f_{q}^{L_{1}}$. We estimate its numerical value $u$ by minimizing the sum $S$ of the squared residues
@@
computed by comparing the low-accuracy eigenvalue estimations with the effect of the high-accuracy constraints from MLCB. This procedure can be readily adjusted to tackle scenarios where different pairs of layers have been studied with MLCB.
Appendix F Random noise models
@@
Inactive parameters are drawn from a Gaussian distribution with mean $m_{w}^{i}$ and standard deviation $\sigma_{w}^{i}$. The active ones are sampled from a Gaussian with mean $m_{w}^{a}$ and standard deviation $\sigma_{w}^{a}$, where, for every gate in the layer, $m_{w}^{a}$ itself is drawn from a Gaussian with mean $\bar{m}_{w}^{a}$ and standard deviation $\bar{\sigma}_{w}^{a}$ . This additional sampling step introduces variability between gates within a single layer, resulting in better and worse-performing ones.
@@
$\begin{split}m_{1}^{i}=2\,10^{-4}\quad&\sigma_{1}^{i}=8\,10^{-4}\\ \bar{m}_{1}^{a}=1\,10^{-3}\quad&\bar{\sigma}_{1}^{a}=7.5\,10^{-4}\\ \quad&\sigma_{1}^{a}=1\,10^{-3}\end{split}$ (43)
@@
$\begin{split}m_{2}^{i}=1.5\,10^{-4}\quad&\sigma_{2}^{i}=1\,10^{-3}\\ \bar{m}_{2}^{a}=2\,10^{-3}\quad&\bar{\sigma}_{2}^{a}=1.5\,10^{-3}\\ \quad&\sigma_{2}^{a}=2\,10^{-3}\end{split}$ (44)
@@
$f^{B}_{XIII}$ $f^{B}_{XIXI}f^{B}_{XIXZ}$ +0.50+0.50 $\begin{pmatrix}-0.50\\ \end{pmatrix}$ $\begin{pmatrix}f^{B}_{IIXZ}f^{B}_{IIXI}\\ \end{pmatrix}$
@@
$f^{G}_{IIXI}$ $f^{G}_{XZXZ}f^{G}_{XZXI}$ 0.50-0.50 $\begin{pmatrix}+0.50\\ -0.50\\ +0.25\\ +0.75\\ \end{pmatrix}$ $\begin{pmatrix}f^{G}_{XZII}f^{G}_{XZII}\\ f^{G}_{IZII}f^{G}_{IZII}\\ f^{G}_{IIXZ}f^{G}_{IZXZ}\\ f^{G}_{IZXI}f^{G}_{IIXI}\\ \end{pmatrix}$
@@
$f^{B}_{XIII}$ $f^{B}_{XIXI}f^{B}_{XIXZ}$ +0.50+0.50 $\begin{pmatrix}+0.25\\ -0.50\\ +0.25\\ -0.25\\ \end{pmatrix}$ $\begin{pmatrix}f^{B}_{IIIZ}f^{B}_{IIIZ}\\ f^{B}_{IIXZ}f^{B}_{IIXI}\\ f^{B}_{XZII}f^{B}_{XIII}\\ f^{B}_{XIIZ}f^{B}_{XZIZ}\\ \end{pmatrix}$
@@
$f^{G}_{IIXI}$ $f^{G}_{XZXZ}f^{G}_{XZXI}$ 0.50-0.50 $\begin{pmatrix}-0.50\\ +0.25\\ +0.75\\ -0.25\\ +0.50\\ -0.25\\ +0.25\\ \end{pmatrix}$ $\begin{pmatrix}f^{G}_{IZII}f^{G}_{IZII}\\ f^{G}_{IIXZ}f^{G}_{IZXZ}\\ f^{G}_{IIXI}f^{G}_{IZXI}\\ f^{G}_{IIIZ}f^{G}_{IIIZ}\\ f^{G}_{XZII}f^{G}_{XZII}\\ f^{G}_{XIII}f^{G}_{XIII}\\ f^{G}_{XIIZ}f^{G}_{XIIZ}\\ \end{pmatrix}$
@@
$f^{B}_{IXII}$ $f^{B}_{IXIX}f^{B}_{ZXIX}$ +0.50+0.50 $\begin{pmatrix}-0.25\\ +0.25\\ -0.50\\ \end{pmatrix}$ $\begin{pmatrix}f^{B}_{ZXII}f^{B}_{ZXZI}\\ f^{B}_{IXZI}f^{B}_{IXII}\\ f^{B}_{IIIX}f^{B}_{IIIX}\\ \end{pmatrix}$
@@
$f^{G}_{IIXI}$ $f^{G}_{XZXZ}f^{G}_{XZXI}$ 0.50-0.50 $\begin{pmatrix}-0.50\\ +0.50\\ +1.00\\ \end{pmatrix}$ $\begin{pmatrix}f^{G}_{IIZI}f^{G}_{IIZI}\\ f^{G}_{IXZI}f^{G}_{ZXZI}\\ f^{G}_{IIZX}f^{G}_{IIIX}\\ \end{pmatrix}$
@@
$f^{B}_{IXII}$ $f^{B}_{IXIX}f^{B}_{ZXIX}$ +0.50+0.50 $\begin{pmatrix}+0.25\\ -0.25\\ -0.25\\ +0.25\\ -0.25\\ \end{pmatrix}$ $\begin{pmatrix}f^{B}_{ZIII}f^{B}_{ZIII}\\ f^{B}_{IIIX}f^{B}_{IIIX}\\ f^{B}_{ZXII}f^{B}_{ZXZI}\\ f^{B}_{IXZI}f^{B}_{IXII}\\ f^{B}_{ZIIX}f^{B}_{ZIIX}\\ \end{pmatrix}$
@@
$f^{G}_{IIXI}$ $f^{G}_{XZXZ}f^{G}_{XZXI}$ 0.50-0.50 $\begin{pmatrix}-0.25\\ +0.25\\ +0.50\\ -0.50\\ +0.75\\ \end{pmatrix}$ $\begin{pmatrix}f^{G}_{ZIII}f^{G}_{ZIII}\\ f^{G}_{ZIIX}f^{G}_{ZIZX}\\ f^{G}_{IXZI}f^{G}_{ZXZI}\\ f^{G}_{IIZI}f^{G}_{IIZI}\\ f^{G}_{IIZX}f^{G}_{IIIX}\\ \end{pmatrix}$